# Learning to Defer in Content Moderation: The Human-AI Interplay

### v2

@@ D.2 Regret guarantee of COLBACID (Theorem 4)
where the first inequality is because $\sum_{k}A_{k}^{\textsc{TABACID}}(t)\leq 1,\sum_{k}\bar{A}_{k}(t)\leq 1$ and $\tilde{Q}_{g(k)}(t)\leq 2\beta r_{\max}\ell_{\max}$ by Lemma 5.6; the last inequality is by Lemma 5.4 that $\mathbb{P}\{\mathcal{E}^{c}\}\leq\delta$ and the assumption that $\delta\leq 1/T$.
We next upper bound the first term in (28). Conditioning on $\mathcal{E}$, we bound $(A_{k}^{\textsc{TABACID}}(t)-\bar{A}_{k}(t))(\beta r_{k}\ell_{k}-\tilde{Q}_{g(k)}(t))$ for a fixed period $t$ and a fixed type $k$. Since $\boldsymbol{\theta}^{\star}\in\mathcal{C}_{t-1}$ by $\mathcal{E}$, we have that $h_{k}$ is inside the confidence interval $[\underline{h}_{k}(t),\bar{h}_{k}(t)]$ and thus $\bar{r}_{k}(t)\geq r_{k}$. If there is no type-$k$ arrival, i.e., $\Lambda_{k}(t)=0$, we have that $A_{k}^{\textsc{TABACID}}(t)=\bar{A}_{k}(t)=0$ and thus the term is zero. Otherwise, since $\bar{r}_{k}(t)\geq r_{k}$, $\bar{A}_{k}(t)=\mathbbm{1}\left(\beta\bar{r}_{k}(t)\ell_{k}\geq\tilde{Q}_{g(k)}(t)\right)\geq\mathbbm{1}\left(\beta r_{k}\ell_{k}\geq\tilde{Q}_{g(k)}(t)\right)=A_{k}^{\textsc{TABACID}}(t)$. Therefore, $(A_{k}^{\textsc{TABACID}}(t)-\bar{A}_{k}(t))(\beta r_{k}\ell_{k}-\tilde{Q}_{g(k)}(t))$ is positive only if $A_{k}^{\textsc{TABACID}}(t)=0$ and $\bar{A}_{k}(t)=1$, under which we have
@@
where the last inequality is because the optimistic admission rule only admits a post ($\bar{A}_{k}(t)=1$) when $\tilde{Q}_{g(k)}(t)\leq\beta\bar{r}_{k}(t)\ell_{k}$. To bound the right hand side of (29), we first bound $\bar{r}_{k}(t)-r_{k}$ by
$\bar{r}_{k}(t)-r_{k}$ $=\max_{h\in[\underline{h}_{k}(t),\bar{h}_{k}(t)]}r_{k}(h)-r_{k}(h_{k})=\max_{\boldsymbol{\theta}\in\mathcal{C}_{t-1}}(r_{k}(\boldsymbol{\phi}_{k}^{\intercal}\boldsymbol{\theta})-r_{k}(\boldsymbol{\phi}_{k}^{\intercal}\boldsymbol{\theta}^{\star}))$
$\leq 1.5v_{k}\max_{\boldsymbol{\theta}\in\mathcal{C}_{t-1}}|\boldsymbol{\phi}_{k}^{\intercal}\boldsymbol{\theta}-\boldsymbol{\phi}_{k}^{\intercal}\boldsymbol{\theta}^{\star}|$
$\leq 1.5v_{k}\max_{\boldsymbol{\theta}\in\mathcal{C}_{t-1}}\left(|\boldsymbol{\phi}_{k}^{\intercal}\boldsymbol{\theta}-\boldsymbol{\phi}_{k}^{\intercal}\hat{\boldsymbol{\theta}}_{t-1}|+|\boldsymbol{\phi}_{k}^{\intercal}\hat{\boldsymbol{\theta}}_{t-1}-\boldsymbol{\phi}_{k}^{\intercal}\boldsymbol{\theta}^{\star}|\right)$
$\leq 1.5v_{k}\max_{\boldsymbol{\theta}\in\mathcal{C}_{t-1}}\left(\|\boldsymbol{\phi}_{k}\|_{\bar{V}_{t-1}^{-1}}\|\boldsymbol{\theta}-\hat{\boldsymbol{\theta}}_{t-1}\|_{\bar{V}_{t-1}}+\|\boldsymbol{\phi}_{k}\|_{\bar{V}_{t-1}^{-1}}\|\hat{\boldsymbol{\theta}}_{t-1}-\boldsymbol{\theta}^{\star}\|_{\bar{V}_{t-1}}\right)$
$\leq 3r_{\max}B_{\delta}(T)\|\boldsymbol{\phi}_{k}\|_{\bar{V}_{t-1}^{-1}}$ (30)
where the last inequality is because we condition on $\mathcal{E}$, which requires $\boldsymbol{\theta}^{\star}\in\mathcal{C}_{t-1}$, the definition of $\mathcal{C}_{t-1}$ in (14), and our definition of $r_{\max}=v_{\max}(H+\sigma_{\max})>v_{\max}$. We complete the proof by
@@
$\leq 3\beta r_{\max}\ell_{\max}B_{\delta}(T)\left(\mathbb{E}\left[\sum_{t=1}^{T}\sum_{k\in\mathcal{K}}\|\boldsymbol{\phi}_{k}\|_{\bar{V}_{t-1}^{-1}}\bar{A}_{k}(t)\right]+1\right)$ (By (30))
$\leq 3\beta r_{\max}\ell_{\max}B_{\delta}(T)\left(d\ln(1+T/d)\left(4GQ_{\max}+\frac{9B_{\delta}^{2}(T)}{\max(\eta,\gamma)^{2}}\right)+\sqrt{2GTd\ln(1+T/d)}\right)$
@@
$\leq 3GQ_{\max}d\ln(1+T/d)+\sqrt{2GTd\ln(1+T/d)}+\frac{9B_{\delta}^{2}(T)d\ln(1+T/d)}{\max(\eta,\gamma)^{2}}$ (By Lemmas 5.5 and D.6)
$\leq d\ln(1+T/d)\left(3GQ_{\max}+\frac{9B_{\delta}^{2}(T)}{\max(\eta,\gamma)^{2}}\right)+\sqrt{2GTd\ln(1+T/d)}.$
Appendix E Some Useful Known Results
@@ E.1 Concentration Inequality
We also use the Chernoff bound for normal random variables.
@@
Given $N$ i.i.d. random variables $X_{i}$ that are normally distributed with variance $\sigma^{2}$. Let $X=\sum_{n=1}^{N}X_{n}$. Then for any $x>0$,
@@
Proof.
We know that $X$ is normally distributed with mean $\mathbb{E}\left[X\right]$ and variance $N\sigma^{2}$. The result then follows from a Chernoff bound on $X$; see [BLM13, page 22]. ∎
@@ E.2 Facts on Matrix Norm
Assume that $\|\boldsymbol{\theta}^{\star}\|_{2},~{}\|\boldsymbol{X}_{t}\|_{2}\leq U$ and that $\eta_{t}$ is conditionally $R-$sub-Gaussian for some $R\geq 0$ such that $\forall u\in\mathbb{R}$, $\mathbb{E}\left[e^{u\eta_{t}}\,\middle|\,\mathcal{F}_{t-1}\right]\leq\exp(u^{2}R^{2}/2)$. Then for any $\delta>0$, with probability at least $1-\delta$, for all $t\geq 0$, $\boldsymbol{\theta}_{\star}$ lies in the set
@@ E.3 Additional Analytical Facts
For any $x$, we have $\varphi^{\prime}(x)-x\varphi(x)$ and $|\varphi^{\prime}(x)|\leq 1/4$.
@@
By definition, $\varphi(x)=\frac{1}{\sqrt{2\pi}}e^{-x^{2}/2}$ and thus $\varphi^{\prime}(x)=\frac{-x}{\sqrt{2\pi}}e^{-x^{2}/2}=-x\varphi(x)$. To bound $|\varphi^{\prime}(x)|$, it suffices to bound it for $x>0$ due to symmetry. Let $f(x)=xe^{-x^{2}/2}$. We have $f^{\prime}(x)=(1-x^{2})e^{-x^{2}/2}$. Then the only $x\geq 0$ with $f^{\prime}(x)=0$ is $x=1$. We know $f(0)=0,f(1)=e^{-1/2}$ and $f(+\infty)=0$. As a result, the maximum value of $f(x)$ is $e^{-1/2}$, which shows $|\varphi^{\prime}(x)|\leq\frac{1}{\sqrt{2\pi}}e^{-1/2}\leq\frac{1}{4}$. ∎
@@
We have that $r_{k}^{O}(h)$ is increasing for $h\leq 0$ and $r_{k}^{R}(h)$ is decreasing for $h\geq 0$. Further,for any $t$ and $k$, we have $\bar{r}_{k}(t)\leq v_{k}\sigma_{k}$.
@@
Since $r_{k}^{O}(h)=r_{k}^{R}(-h)$ for $h<0$ by (5), we only need to show $r_{k}^{R}(h)$ is decreasing for $h\geq 0$. We know that $r_{k}^{R}(h)=v_{k}(-h\Phi(-h/\sigma_{k})+\sigma_{k}\varphi(-h/\sigma_{k}))$. Then its derivative is given by
$v_{k}\left(-\Phi(-h/\sigma_{k})+\frac{h}{\sigma_{k}}\varphi(-h/\sigma_{k})-\frac{h}{\sigma_{k}}\varphi(-h/\sigma_{k})\right)=-v_{k}\Phi(-h/\sigma_{k})<0.$
As a result, $r_{k}^{R}(h)$ is decreasing for $h\geq 0$, and is maximized at $h=0$, which also shows $r_{k}^{O}(h)$ is increasing for $h<0$.
By definition, $\bar{r}_{k}(t)=\max_{h\in[\underline{h}_{k}(t),\bar{h}_{k}(t)]}r_{k}(h)\leq\max_{h}r_{k}(h)=r_{k}(0)=r_{k}^{R}(0)$. We finish the proof by noting that $r_{k}^{R}(0)=v_{k}\sigma_{k}$. ∎

### v3

@@ D.2 Regret guarantee of COLBACID (Theorem 4)
We first need a lemma that bounds $\bar{r}_{k}(t)-r_{k}$.
Lemma D.5.
Conditioned on $\mathcal{E}$, we have $\bar{r}_{k}(t)-r_{k}\leq 2B_{\delta}(T)\|\boldsymbol{\phi}_{k}\|_{\bar{V}_{t-1}^{-1}}$ for any type $k$.
Proof.
Conditioning on $\mathcal{E}$ gives $\boldsymbol{\theta}^{\star,O}\in\mathcal{C}^{O}_{t-1}$ and $\boldsymbol{\theta}^{\star,O}\in\mathcal{C}^{R}_{t-1}$. As a result, for any $\boldsymbol{\theta}^{O}\in\mathcal{C}^{O}_{t-1}$,
$|\boldsymbol{\phi}_{k}^{\intercal}\boldsymbol{\theta}^{O}-r_{k}^{O}|=|\boldsymbol{\phi}_{k}^{\intercal}\boldsymbol{\theta}^{O}-\boldsymbol{\phi}_{k}^{\intercal}\boldsymbol{\theta}^{\star,O}|$ $\leq|\boldsymbol{\phi}_{k}^{\intercal}\boldsymbol{\theta}^{O}-\boldsymbol{\phi}_{k}^{\intercal}\hat{\boldsymbol{\theta}}^{O}(t-1)|+|\boldsymbol{\phi}_{k}^{\intercal}\hat{\boldsymbol{\theta}}^{O}(t-1)-\boldsymbol{\phi}_{k}^{\intercal}\boldsymbol{\theta}^{\star,O}|$
$\leq\|\boldsymbol{\phi}_{k}\|_{\bar{V}_{t-1}^{-1}}\left(\|\boldsymbol{\theta}^{O}-\hat{\boldsymbol{\theta}}^{O}(t-1)\|_{\bar{V}_{t-1}}+\|\boldsymbol{\theta}^{\star,O}-\hat{\boldsymbol{\theta}}^{O}(t-1)\|_{\bar{V}_{t-1}}\right)$ (Cauchy-Schwarz inequality)
$\leq 2B_{\delta}(T)\|\boldsymbol{\phi}_{k}\|_{\bar{V}_{t-1}^{-1}}.$ (By (15) and that $\boldsymbol{\theta}^{\star,O},\boldsymbol{\theta}^{O}\in\mathcal{C}^{O}_{t-1}$)
Similarly, we also have $|\boldsymbol{\phi}_{k}^{\intercal}\boldsymbol{\theta}^{R}-r_{k}^{R}|\leq 2B_{\delta}(T)\|\boldsymbol{\phi}_{k}\|_{\bar{V}_{t-1}^{-1}}$ for any $\boldsymbol{\theta}^{R}\in\mathcal{C}_{t-1}^{R}.$ As a result, for any $\boldsymbol{\theta}^{O}\in\mathcal{C}_{t-1}^{O},\boldsymbol{\theta}^{R}\in\mathcal{C}_{t-1}^{R}$,
$\min(\boldsymbol{\phi}_{k}^{\intercal}\boldsymbol{\theta}^{O},\boldsymbol{\phi}_{k}^{\intercal}\boldsymbol{\theta}^{R})\leq\min(r_{k}^{O}+2B_{\delta}(T)\|\boldsymbol{\phi}_{k}\|_{\bar{V}_{t-1}^{-1}},r_{k}^{R}+2B_{\delta}(T)\|\boldsymbol{\phi}_{k}\|_{\bar{V}_{t-1}^{-1}})=\min(r_{k}^{O},r_{k}^{R})+2B_{\delta}(T)\|\boldsymbol{\phi}_{k}\|_{\bar{V}_{t-1}^{-1}}.$
The definition of $\bar{r}_{k}(t)$ in (17) then gives
$\bar{r}_{k}(t)-r_{k}=\max_{\boldsymbol{\theta}^{O}\in\mathcal{C}_{t-1}^{O},\boldsymbol{\theta}^{R}\in\mathcal{C}_{t-1}^{R}}\min(\boldsymbol{\phi}_{k}^{\intercal}\boldsymbol{\theta}^{O},\boldsymbol{\phi}_{k}^{\intercal}\boldsymbol{\theta}^{R})-\min(r_{k}^{O},r_{k}^{R})\leq 2B_{\delta}(T)\|\boldsymbol{\phi}_{k}\|_{\bar{V}_{t-1}^{-1}}.$
@@
where the first inequality is by $\sum_{k}A_{k}^{\textsc{TABACID}}(t)\leq 1,\sum_{k}\bar{A}_{k}(t)\leq 1$ and $\tilde{Q}_{g(k)}(t)\leq 2\beta r_{\max}\ell_{\max}$ by Lemma 5.6; the last one is by Lemma 5.4 that $\mathbb{P}\{\mathcal{E}^{c}\}\leq 2\delta$ and the assumption that $\delta\leq 1/(2T)$.
We next upper bound the first term in (29). Conditioning on $\mathcal{E}$, we bound $(A_{k}^{\textsc{TABACID}}(t)-\bar{A}_{k}(t))(\beta r_{k}\ell_{k}-\tilde{Q}_{g(k)}(t))$ for a fixed period $t$ and a fixed type $k$. Since $\boldsymbol{\Theta}^{\star}\in\mathcal{C}_{t-1}$ by $\mathcal{E}$, we have that $\bar{r}_{k}(t)\geq r_{k}$. If there is no type-$k$ arrival, i.e., $\Lambda_{k}(t)=0$, we have that $A_{k}^{\textsc{TABACID}}(t)=\bar{A}_{k}(t)=0$ and thus the term is zero. Otherwise, since $\bar{r}_{k}(t)\geq r_{k}$, $\bar{A}_{k}(t)=\mathbbm{1}\left(\beta\bar{r}_{k}(t)\ell_{k}\geq\tilde{Q}_{g(k)}(t)\right)\geq\mathbbm{1}\left(\beta r_{k}\ell_{k}\geq\tilde{Q}_{g(k)}(t)\right)=A_{k}^{\textsc{TABACID}}(t)$. Therefore, $(A_{k}^{\textsc{TABACID}}(t)-\bar{A}_{k}(t))(\beta r_{k}\ell_{k}-\tilde{Q}_{g(k)}(t))$ is positive only if $A_{k}^{\textsc{TABACID}}(t)=0$ and $\bar{A}_{k}(t)=1$, under which we have
@@
where the last inequality is because the optimistic admission rule only admits a post ($\bar{A}_{k}(t)=1$) when $\tilde{Q}_{g(k)}(t)\leq\beta\bar{r}_{k}(t)\ell_{k}$. We complete the proof by
@@
$\leq 3\beta r_{\max}\ell_{\max}B_{\delta}(T)\left(\mathbb{E}\left[\sum_{t=1}^{T}\sum_{k\in\mathcal{K}}\|\boldsymbol{\phi}_{k}\|_{\bar{V}_{t-1}^{-1}}\bar{A}_{k}(t)\right]+1\right)$ (By Lemma D.5)
$\leq 3\beta r_{\max}\ell_{\max}B_{\delta}(T)\left(d\ln(1+T/d)\left(4GQ_{\max}+\frac{34B_{\delta}^{2}(T)}{\max(\eta,\gamma)^{2}}\right)+\sqrt{2GTd\ln(1+T/d)}\right)$
@@
$\leq 3GQ_{\max}d\ln(1+T/d)+\sqrt{2GTd\ln(1+T/d)}+\frac{34B_{\delta}^{2}(T)d\ln(1+T/d)}{\max(\eta,\gamma)^{2}}$ (By Lemmas 5.5 and D.7)
$\leq d\ln(1+T/d)\left(3GQ_{\max}+\frac{34B_{\delta}^{2}(T)}{\max(\eta,\gamma)^{2}}\right)+\sqrt{2GTd\ln(1+T/d)}.$
Appendix E Analytical Facts
@@ E.1 Concentration Inequality
We also use the Chernoff bound for sub-Gaussian random variables, which is given in [BLM13, Page 25].
@@
Given $n$ i.i.d. random variables $X_{i}$ that are sub-Gaussian with variance proxy $\sigma^{2}$ and letting $X=\sum_{i=1}^{n}X_{i}/n$, for any $x>0$,
@@ E.2 Facts on Matrix Norm
Assume that $\|\boldsymbol{\theta}^{\star}\|_{2},~{}\|\boldsymbol{X}_{t}\|_{2}\leq U$ and that $\eta_{t}$ is conditionally sub-Gaussian with variance proxy $R^{2}$ such that $\forall u\in\mathbb{R}$, $\mathbb{E}\left[e^{u\eta_{t}}\,\middle|\,\mathcal{F}_{t-1}\right]\leq\exp(u^{2}R^{2}/2)$. Then for any $\delta>0$, with probability at least $1-\delta$, for all $t\geq 0$, $\boldsymbol{\theta}_{\star}$ lies in the set
@@ E.3 Additional Analytical Facts
If $X$ is a sub-Gaussian random variable with variance proxy $\sigma^{2}$, then $-X$ is sub-Gaussian with variance proxy $\sigma^{2}$.
@@
For any $s\in\mathbb{R}$, $\mathbb{E}\left[\exp(s(-X-\mathbb{E}\left[-X\right]))\right]=\mathbb{E}\left[\exp((-s)(X-\mathbb{E}\left[X\right])\right]\leq\exp(s^{2}\sigma^{2}/2)$. As a result, $-X$ is sub-Gaussian with variance proxy $\sigma^{2}$. ∎
@@
Given two independent sub-Gaussian random variables $X,Y$ with variance proxy $\sigma_{X}^{2}$ and $\sigma_{Y}^{2}$ respectively, their sum $X+Y$ is sub-Gaussian with variance proxy $\sigma_{X}^{2}+\sigma_{Y}^{2}.$
@@
For any $s\in\mathbb{R}$, $\mathbb{E}\left[\exp(s(X+Y-\mathbb{E}\left[X+Y\right]))\right]=\mathbb{E}\left[\exp(s(X-\mathbb{E}\left[X\right])\right]\mathbb{E}\left[\exp(s(Y-\mathbb{E}\left[Y\right])\right]\leq\exp(s^{2}(\sigma_{X}^{2}+\sigma_{Y}^{2})/2)$, where the equality uses independence between $X,Y$. ∎